OpenStudy (anonymous):

(1/3)^x=1/9 solve the equation for x. Round to the nearest two decimal places

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

\[3^{-x} = 3^{-2}\] -x = -2 x = 2

OpenStudy (anonymous):

ok thanks to both of you can you tell me if im correct when I say e^x=4 and x=1.39

OpenStudy (anonymous):

take logarithm to the base e, and check your log book for Log 4 base e..

OpenStudy (anonymous):

(1/3)^x = (1/9) Solution: => ln(1/3)^x = ln(1/9) {multiply both sides by natural log ln} => x*ln(1/3) = ln(1/9) {exponent rule for log/ln} => x = ln(1/9) / ln(1/3) => x = 2

OpenStudy (anonymous):

Pepsa, you are correct.